### v1

# Schwarzschild-Levi-Civita black hole

Abstract
Inspired by the geometry of the Ernst black hole which interpolates the static spherically symmetric Schwarzschild black hole and the static axially symmetric magnetic Melvin spacetime, in this Letter, we aim to present a solution to Einstein’s field equation in vacuum to interpolate the Schwarzschild black hole and the Levi-Civita spacetime. Hence, we shall call the solution the Schwarzschild-Levi-Civita black hole. Effectively the solution is a one-parameter non-asymptotically flat black hole that is static, axially symmetric, and singular in its axis of symmetry.
Schwarszchild black hole; Leci-Civita metric; Vacuum black hole; Exact solution;
I Introduction
One of the axially symmetric black hole spacetime is the so-called Ernst black hole Ernst with the line element
$ds^{2}=\Lambda^{2}\left(-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{1-\frac{2m}{r}}+r^{2}d\theta^{2}\right)+\frac{r^{2}\sin^{2}\theta d\phi^{2}}{\Lambda^{2}},$ (1)
in which
$\Lambda=1+\frac{1}{4}B^{2}r^{2}\sin^{2}\theta.$ (2)
The black hole (1) is a 2-parameter static axially symmetric solution of Einstein-Maxwell equations with an external magnetic field $B$ in the polar direction. The solution reduces into the Schwarzschild black hole in the absence of the magnetic field i.e., $B=0,$ and coincides with Melvin magnetic universe with $m=0$. For these two extreme limits usually the solution (1) is also referred to as the Schwarzschild black hole immersed into a magnetic field or the magnetized Schwarzschild black hole. The Ernst black hole has been widely studied in the literature. One of the very recent studies was done by Senjaya in Senjaya where the exact massless scalar quasibound state in the Ernst black hole background is presented. In Becar Becar et al. studied the quasinormal modes of a charged scalar field in the same background spacetime. Gibbons et al. in Gibbons reconsidered the original rotating Ernst black hole and showed that it asymptotically doesn’t mimic the Melvin metric unless the electric charge of the black hole and the external magnetic field satisfy $Q=jB\left(1+\frac{1}{4}j^{2}B^{4}\right)$ in which $Q$ and $B$ are the electric charge of the black hole and the external magnetic field, respectively, and $j$ is the angular momentum of the Kerr-Newman black hole which was used as a seed solution in the Ernst’ procedure Wild . Furthermore, there are significant details on Ernst’s procedure and the rotating Ernst solution with the electromagnetic fields in Gibbons which makes it essential in studying Ernst’s black hole.
In this current study, we present a new black hole solution that is similar to the Ernst static black hole (1) but instead of interpolating the Schwarzschild black hole and Melvin magnetic universe, it interpolates the Schwarzschild black hole and the Levi-Civita spacetime. Both the Schwarzschild black hole and the Levi-Civita spacetime Levi are solutions to Einstein’s equations in vacuum. Hence, unlike the Ernst black hole, the Schwarzschild-Levi-Civita black hole - this is what we shall call our solution hereafter - is also a solution of Einstein’s field equation in vacuum. In general relativity, except the Schwarzschild black hole and the Levi-Civita spacetime which are static, there exists another solution to Einstein’s equations in vacuum which is known as $\gamma$-metric or Zipoy-Voorhees spacetime Z ; V . All these solutions are very well-known in the literature due to their exactness and clarity which make them practical. Having an exact solution to Einstein’s field equation in its own right is very important and deserves to be searched for. In the solution we present in this Letter, we observe not only its exactness but also its simplicity and clarity which in our opinion can be of interest and significant in the community of general relativity.
II Field equations and the black hole solution
Inspired by the Ernst black hole solution (1), we consider the Schwarzschild-Levi-Civita black hole to be described by the line element
$ds^{2}=\Lambda^{2}\left(-f\left(r\right)dt^{2}+\frac{dr^{2}}{f\left(r\right)}+r^{2}d\theta^{2}\right)+\Lambda^{-2}r^{2}\sin^{2}\theta d\phi^{2},$ (3)
in which in (3) $\Lambda=\Lambda\left(r,\theta\right)$. We aim to solve Einstein’s vacuum field equations i.e., $G_{\mu}^{\nu}=0$. Concerning the static line element (3) one explicitly obtains the field equations given by
$G_{t}^{t}=\frac{f\Lambda_{,r}^{2}}{\Lambda^{4}}-\frac{2f\Lambda_{,r}}{r\Lambda^{3}}+\frac{f^{\prime}}{r\Lambda^{2}}+\frac{\Lambda_{,\theta}^{2}}{r^{2}\Lambda^{4}}-\frac{2\Lambda_{,\theta}}{r^{2}\Lambda^{3}}\cot\theta-\frac{1-f}{r^{2}\Lambda^{2}}=0,$ (4)
$G_{r}^{r}=-\frac{f\Lambda_{,r}^{2}}{\Lambda^{4}}+\frac{2f\Lambda_{,r}}{r\Lambda^{3}}+\frac{f^{\prime}}{r\Lambda^{2}}+\frac{\Lambda_{,\theta}^{2}}{r^{2}\Lambda^{4}}-\frac{2\Lambda_{,\theta}}{r^{2}\Lambda^{3}}\cot\theta-\frac{1-f}{r^{2}\Lambda^{2}}=0,$ (5)
$G_{r}^{\theta}=\frac{2\Lambda_{,r}\Lambda_{,\theta}}{r^{2}\Lambda^{4}}+\frac{2\Lambda_{,r}}{r^{2}\Lambda^{3}}\cot\theta+\frac{2\Lambda_{,\theta}}{r^{2}\Lambda^{3}}=0,$ (6)
$G_{\theta}^{\theta}=\frac{f\Lambda_{,r}^{2}}{\Lambda^{4}}-\frac{2f\Lambda_{,r}}{r\Lambda^{3}}+\frac{f^{\prime}}{r\Lambda^{2}}+\frac{\Lambda_{,\theta}^{2}}{r^{2}\Lambda^{4}}-\frac{2\Lambda_{,\theta}}{r^{2}\Lambda^{3}}\cot\theta+\frac{f^{\prime\prime}}{2\Lambda^{2}}=0,$ (7)
and
$G_{\phi}^{\phi}=\frac{2f\Lambda_{,r}}{r\Lambda^{3}}+\frac{2f^{\prime}\Lambda_{,r}}{\Lambda^{3}}+\frac{2f\Lambda_{,rr}}{\Lambda^{3}}+\frac{2\Lambda_{,\theta\theta}}{r^{2}\Lambda^{3}}-\frac{\Lambda_{,\theta}^{2}}{r^{2}\Lambda^{4}}-\frac{f\Lambda_{,r}^{2}}{\Lambda^{4}}+\frac{f^{\prime}}{r\Lambda^{2}}+\frac{f^{\prime\prime}}{2\Lambda^{2}}=0,$ (8)
where a prime stands for the derivative with respect to $r$ and $\Lambda_{,x}=\frac{\partial\Lambda}{\partial x}.$ Combining (4) and (5) yields
$\Lambda_{,r}\left(r\Lambda_{,r}-2\Lambda\right)=0,$ (9)
which admits two solutions, given by
$\Lambda\left(r,\theta\right)=\lambda\left(\theta\right),$ (10)
and
$\Lambda\left(r,\theta\right)=\lambda\left(\theta\right)r^{2},$ (11)
in which $\lambda\left(\theta\right)$ is an arbitrary function of $\theta.$ From (6), we observe that the first solution i.e., Eq. (10) yields the trivial solution $\lambda\left(\theta\right)=const.$ which leads to the Schwarzschild black hole. On the other hand $\Lambda\left(r,\theta\right)=\lambda\left(\theta\right)r^{2}$ is the non-trivial solution which after introducing into (6) implies
$\lambda^{\prime}\left(\theta\right)-2\cot\theta\lambda\left(\theta\right)=0.$ (12)
The latter equation is solved to obtain the exact form of $\lambda\left(\theta\right)$ which is given by
$\lambda\left(\theta\right)=\lambda_{0}\sin^{2}\theta,$ (13)
where $\lambda_{0}$ is an integration constant. Upon considering (13) and (11) in the field equations (4)-(8) they reduce to the following equations
$G_{t}^{t}=G_{r}^{r}=\frac{rf^{\prime}+f-1}{\lambda_{0}^{2}r^{6}\sin^{4}\theta}=0,$ (14)
$G_{\theta}^{\theta}=\frac{rf^{\prime\prime}+2f^{\prime}}{2\lambda_{0}^{2}r^{5}\sin^{4}\theta}=0,$ (15)
and
$G_{\phi}^{\phi}=\frac{r^{2}f^{\prime\prime}+10rf^{\prime}+8f-8}{\lambda_{0}^{2}r^{6}\sin^{4}\theta}=0.$ (16)
Upon solving (14) which satisfies the other two equations too, the solution for $f\left(r\right)$ is found to be
$f\left(r\right)=1-\frac{2m}{r},$ (17)
where $m$ is an integration constant related to the mass of the black hole. Finally, in the spherically symmetric coordinates i.e., $\left\{t,r,\theta,\phi\right\},$ the black hole solution is expressed by the line element
$ds^{2}=\lambda_{0}^{2}r^{4}\sin^{4}\theta\left(-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{1-\frac{2m}{r}}+r^{2}d\theta^{2}\right)+\frac{d\phi^{2}}{\lambda_{0}^{2}r^{2}\sin^{2}\theta}.$ (18)
This is the solution of Einstein’s vacuum field equations i.e., $G_{\mu\nu}=0,$ and since the Ricci scalar is also zero they are the solution of $R_{\mu\nu}=0$, too. Although the Ricci scalar and the Ricci tensors are zero the Kretschmann scalar is non-zero and is given by
$K=\frac{1008m^{2}}{\lambda_{0}^{4}r^{14}\sin^{8}\theta}-\frac{288\left(2+\sin^{2}\theta\right)m}{\lambda_{0}^{4}r^{13}\sin^{10}\theta}+\frac{192}{\lambda_{0}^{4}r^{12}\sin^{12}\theta}.$ (19)
For convenience one may express (18) in the prolate coordinates i.e., $\left\{T,X,Y,\phi\right\}$ where $T=\sqrt{\lambda_{0}}t,$ $X=\sqrt{\lambda_{0}}r-1$ and $Y=\cos\theta$. Hence, (18) reads
$d\tilde{s}^{2}=\lambda_{0}ds^{2}=\left(X+1\right)^{4}\left(1-Y^{2}\right)^{2}\left(-\left(1-\frac{2\tilde{m}}{X+1}\right)dT^{2}+\frac{dX^{2}}{1-\frac{2\tilde{m}}{X+1}}+\frac{\left(X+1\right)^{2}}{1-Y^{2}}dY^{2}\right)+\frac{d\phi^{2}}{\left(X+1\right)^{2}\left(1-Y^{2}\right)},$ (20)
in which $\tilde{m}=\sqrt{\lambda_{0}}m.$ We add that while $r\in\left[0,\infty\right)$ and $\theta\in\left[0,\pi\right]$, the prolate coordinates are defined such that $X\in\left[-1,\infty\right)$ and $Y\in\left[-1,1\right].$ In terms of the prolate coordinates the Kretschmann scalar is given by
$K=\frac{1008\tilde{m}^{2}\lambda_{0}^{2}}{\left(X+1\right)^{14}\left(1-Y^{2}\right)^{4}}-\frac{288\left(3-Y^{2}\right)\tilde{m}\lambda_{0}^{2}}{\left(X+1\right)^{13}\left(1-Y^{2}\right)^{5}}+\frac{192\lambda_{0}^{2}}{\left(X+1\right)^{12}\left(1-Y^{2}\right)^{6}}.$ (21)
From (19) and (21), we observe that the axis of symmetry i.e., the line $\theta=0,\pi/Y=\pm 1$ is the singularity of the black hole. Such singularity is due to the mass distribution on the symmetry axis. With $m=0$ the solution (18) can be expressed in cylindrical coordinates i.e., $\left\{t,\rho,z,\phi\right\}$ where $\rho=r\sin\theta$ and $z=r\cos\theta.$ In such setting (18) reads as
$ds^{2}=\lambda_{0}^{2}\rho^{4}\left(-dt^{2}+d\rho^{2}+dz^{2}\right)+\frac{d\phi^{2}}{\lambda_{0}^{2}\rho^{2}}.$ (22)
Remarkably, with $\lambda_{0}=1$ (22) is the well-known Levi-Civita spacetime with the general line element given by
$ds^{2}=\rho^{2\sigma\left(\sigma-1\right)}\left(-dt^{2}+d\rho^{2}\right)+\rho^{2\sigma}dz^{2}+\rho^{2\left(1-\sigma\right)}d\phi^{2},$ (23)
where the Levi-Civita constant is $\sigma=2$. Therefore, the black hole solution (18) may be called the Schwarzschild-Levi-Civita black hole as we mentioned before. The asymptotic behavior of the black hole solution (18) is the same as the Levi-Civita metric i.e.,
$\lim_{r\rightarrow\infty}ds^{2}\rightarrow\lambda_{0}^{2}r^{4}\sin^{4}\theta\left(-dt^{2}+dr^{2}+r^{2}d\theta^{2}\right)+\frac{d\phi^{2}}{\lambda_{0}^{2}r^{2}\sin^{2}\theta},$ (24)
which is the same as Eq. (22). Given these behaviors, we interpret the line element (18) to be the Schwarzschild black hole of mass m in the Levi-Civita spacetime.
III The photon orbit
In this part, we investigate the possible circular geodesic of a null particle. To do so, we start with the effective Lagrangian density of a null particle moving in the vicinity of the black hole (18) that is given by
$2\mathcal{L}=r^{4}\sin^{4}\theta\left(-\left(1-\frac{r_{+}}{r}\right)\dot{t}^{2}+\frac{\dot{r}^{2}}{1-\frac{r_{+}}{r}}+r^{2}\dot{\theta}^{2}\right)+\frac{\dot{\phi}^{2}}{r^{2}\sin^{2}\theta},$ (25)
in which an over-dot stands for the derivative with respect to an affine parameter and $r_{+}=2m$ is the event horizon of the black hole. The conserved energy and angular momentum are given by
$E=-\frac{\partial\mathcal{L}}{\partial\dot{t}}=r^{4}\sin^{4}\theta\left(1-\frac{r_{+}}{r}\right)\dot{t},$ (26)
and
$\ell=\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=\frac{\dot{\phi}}{r^{2}\sin^{2}\theta}.$ (27)
The $\theta$-component of the geodesic equation is not trivially satisfied by a constant $\theta,$ however, it is satisfied by $\theta=\frac{\pi}{2}$. Therefore, for the rest of this section, we set $\theta=\frac{\pi}{2}$ which corresponds to the geodesic of the null particle on the equatorial plane. The main geodesic equation is obtained by applying the constraint $\dot{x}_{\mu}\dot{x}^{\mu}=0$ for the null geodesic which implies
$r^{8}\dot{r}^{2}+r^{6}\ell^{2}\left(1-\frac{r_{+}}{r}\right)=E^{2}.$ (28)
For the circular orbit we set $r=r_{c}$ and consequently $\dot{r}=\ddot{r}=0$ such that considering (28) we find
$r_{c}^{5}\left(r_{c}-r_{+}\right)=\frac{E^{2}}{\ell^{2}},$ (29)
and
$r_{c}=\frac{5}{6}r_{+}.$ (30)
The latter shows that $r_{c}<r_{+}$ and therefore it is inside the black hole. In other words, there is no circular orbit outside the black hole which is unlike the Schwarzschild black hole whose photonsphere located at $r_{c}=\frac{3}{2}r_{+}.$ This simple geodesic implies that the Schwarzschild black hole formed in the Levi-Civita spacetime possesses different fundamental structure than the standard Schwarzschild black hole.
IV Conclusion
In this compact Letter, we introduced a new exact solution to Einstein’s field equation in vacuum. The idea of the existence of such a solution was taken from the Ernst black hole which is an interpolation between the Schwarzschild black hole and the Melvin universe. Our exact solution is effectively a one-parameter black hole that interpolates the Schwarzschild black hole and the Levi-Civita spacetime such that we called it the Schwarzschild-Levi-Civita black hole. It is static and axially symmetric with the singularity at its axis of symmetry. Unlike the Ernst black hole which is the solution of the Einstein-Maxwell field equations the Schwarzschild-Levi-Civita black hole is the solution of the Einstein equations in vacuum. We have studied the existence of the photonsphere and shown that unlike in the Schwarzschild black hole, there exists no photonsphere outside the black hole. This is an indication that this black hole is a non-trivial extension of the Schwarzschild black hole. In this Letter, we haven’t studied the physical properties of the new solution and it is left to a more detailed article in our future work.